# Scaling of Stochastic Normalizing Flows in $\mathrm{SU}(3)$ lattice gauge theory

### v1

@@ II Non-equilibrium Markov Chain Monte Carlo
A non-equilibrium Markov Chain Monte Carlo (NE-MCMC) simulation is composed of an ensemble of out-of-equilibrium evolutions from a prior distribution $q_{0}=\exp(-S_{0})/Z_{0}$ to the target distribution $p=\exp(-S)/Z$ that we wish to sample from. Each of these evolutions is generated by drawing a sequence of $n_{\mathrm{step}}$ configurations $U_{n}$ using a varying transition probability. The starting configuration $U_{0}$ of this sequence is sampled from the prior $q_{0}$ and then driven towards the target distribution $p$ through a sequence of Monte Carlo updates defined by the transition probabilities $P_{c(n)}$. These transition probabilities depend on a set of parameters $c(n)$, known as the protocol, which controls the intermediate steps of the system’s evolution, and drive the configurations sequence from the initial state toward the target distribution, i.e.
@@
where we note that the last transition probability is fixed to be the one of the target distribution, i.e. $P_{c(n_{\mathrm{step}})}\equiv P_{p}$.
Formally, after fixing $n_{\mathrm{step}}$ and the protocol $c(n)$, we can introduce the forward $\mathcal{P}_{\mathrm{f}}$ probability distribution of each evolution $[U_{0},U_{1},\dots,U]\equiv\mathcal{U}$, which can be written as:
@@
similarly, the reverse probability distribution $\mathcal{P}_{\mathrm{r}}$ takes the form of:
@@
In this expression, $\Delta F=\log Z/Z_{0}$ is the free energy difference between the states described by $q_{0}$ and $p$, and $W$ represents the dimensionless work done on the system during its transformation from the initial to the final state:
@@
The computation of free energy differences between two distributions using Monte Carlo simulations is notoriously cumbersome, as it generally requires a large number of samples from intermediate distributions at equilibrium. Thus, eq. (13) represents a powerful estimator of free energy differences, as the intermediate distributions can be sampled without ever letting the system thermalize. In the context of lattice gauge theories Jarzynski’s equality has seen a wide range of applications in the last few years, see refs. [43, 44, 45, 46, 47]. Furthermore, eqs. (13) and  (11) represent the main ingredients of NE-MCMC. i.e. they provide a framework to sample problematic distributions from non-equilibrium distributions in an unbiased way; we refer to ref. [49] for a recent application on a model with frozen topology.
Note that no assumptions regarding thermalization were made during the construction of the NE-MCMC evolutions: the only exception is that if the prior distribution $q_{0}$ is sampled with a MCMC, it must be at equilibrium. While the starting and ending states of the transformations are fixed, there are in principle no restrictions on the protocol, either in the velocity of the evolution (i.e. $n_{\mathrm{step}}$) or in the functional form of $c(n)$; thus, the intermediate and final configurations $U_{n}$, with $0<n\leq n_{\mathrm{step}}$, are in principle allowed to lie arbitrarily far from equilibrium.
@@
It measures the similarity between two distributions and, in this non-equilibrium context, provides a way to study in a quantitative way the reversibility of the stochastic transformation we are employing. Large dissipation, which thermodynamically corresponds to strong irreversibility, leads to a large KL divergence; conversely, when the evolution is fully reversible forward and reverse evolutions are indistinguishable, resulting in $\tilde{D}_{\mathrm{KL}}(q_{0}\mathcal{P}_{\mathrm{f}}\|p\mathcal{P}_{\mathrm{r}})=0$. Intriguingly, since the KL divergence is positive-definite eq. (14) implies the Second Law of thermodynamics for these non-equilibrium transformations, i.e.
@@
The KL divergence of eq. (14) represents a crucial metric to determine the reliability of NE-MCMC when sampling a finite number of evolutions. To understand why, we start by looking at the probability distribution of the last configuration $U\equiv U_{n_{\mathrm{step}}}$, i.e. the one generated by a given protocol at the end of the evolution. We can write it as:
@@
One simple and effective approach to either reduce the KL divergence of eq. (14) or to increase the ESS of eq. (19) is to increase the number of steps in the NE-MCMC protocols, i.e. going towards the $n_{\mathrm{step}}\to\infty$ limit: smoother, “quasi-static” transformations keep the thermodynamic transformations close to equilibrium, thereby maintaining vanishing dissipation. A more elaborate and potentially much more effective strategy is to optimize the functional form of the protocol $c(n)$ itself: this approach remains one of the main challenges for implementing more efficient NE-MCMC algorithms.
@@ III Stochastic Normalizing Flows
$\begin{split}W(\mathcal{U})=\sum_{n=0}^{n_{\mathrm{step}}-1}&S_{c(n+1)}\left[g_{n+1}(U_{n})\right]-S_{c(n)}\left[g_{n+1}(U_{n})\right]+\\ &-\log J_{g_{n+1}}(U_{n}).\end{split}$ (26)
@@ III.1 Gauge Equivariant Layers
The optimal stout-smearing parameters $\rho^{(n)}_{\mu\nu}(x)$ in eq. (33) have to be determined using a training procedure. In this work we simply set them to be invariant under translations and (discrete) rotations, i.e. $\rho^{(n)}_{\mu\nu}(x)=\rho^{(n)}$, for a total of 8 parameters per layer. This construction makes the Jacobian of the transformation easily tractable; for a discussion of invertibility we refer to the discussion on the link-level flow in ref. [29].
@@ IV.1 Training SNFs
The search for the optimal parameters of a SNF architecture is performed through the minimization of the KL divergence of eq. (14). This minimization procedure has a clear thermodynamic interpretation: the parameters of the intermediate stout smearing layers are tuned in such a way that the evolution becomes as reversible as possible, i.e. towards $\langle W_{\mathrm{d}}(\mathcal{U})\rangle_{\mathrm{f}}=0$. It is interesting to note that the structure of the work of eq. (26) is rather peculiar, as it can be expressed as a sum over the steps $n$, each term containing a “stochastic” contribution from the difference of the action and the “deterministic” contribution from the Jacobian of the gauge-equivariant layer. It is then very natural to ask whether the training itself can be performed step by step, i.e. backpropagating the gradients not over the whole flow, but only between two MCMC updates: in practice, one uses the loss
$S_{c(n+1)}\left[g_{n+1}(U_{n})\right]-S_{c(n)}\left[g_{n+1}(U_{n})\right]-\log J_{g_{n+1}}(U_{n})$
@@ IV.2 Scaling of NE-MCMC and SNFs
In this section we compare the performances of NE-MCMC and trained SNFs using the metrics introduced in section II, for different ensembles of transformations in the lattice spacing. In fig. 2 results for the KL divergence of eq. (14) and for the ESS of eq. (19) are presented for two volumes and for several values of $n_{\mathrm{step}}$.
@@
First of all, we observe that when $n_{\mathrm{step}}$ increases, i.e. when evolutions are performed closer to equilibrium, the KL divergence decreases for all architectures, reaching rather quickly values smaller than 1. As expected, the results indicate a convergence to 00 as well, i.e. reversibility appears to be reached asymptotically. Furthermore, if we look at the metrics for each volume separately, it appears that SNFs are roughly a factor 2 more efficient than the standard NE-MCMC: indeed, the same value of $\tilde{D}_{\mathrm{KL}}(q_{0}\mathcal{P}_{\mathrm{f}}\|p\mathcal{P}_{\mathrm{r}})$ is reached for about half the value of $n_{\mathrm{step}}$.
@@
At this point it is important to stress two relevant details. First, that for the same value of $n_{\mathrm{step}}$ SNFs also perform a stout-smearing transformation of the links for each Monte Carlo update; in our experience the increase of the computational cost is however moderate, as in our implementation they are about 25% as expensive as a full MCMC update. Second, this improvement came at a very modest cost: as explained previously, the training was performed uniquely on values of $n_{\mathrm{step}}$ which are much smaller than the ones showed in fig. 2. In practical terms, the computational effort of the training is just a small fraction of the total effort used for the sampling for larger $n_{\mathrm{step}}$.
Now we come to the analysis of the scaling. In fig. 3 we plotted the results for the KL divergence and the ESS as a function of the number of steps normalized by the volume in lattice units. In this way, the results of fig. 2 collapse on a unique curve, along those for $L/a=10$ and $L/a=20$: at least approximately, both metrics can be considered to be a function of $n_{\mathrm{step}}/(L/a)^{4}$. In other words, if we keep the KL divergence or the $\hat{\mathrm{ESS}}$ fixed, i.e. we fix a target for our flow-based sampling, the number of steps $n_{\mathrm{step}}$ needed to reach this target scales with the volume of the lattice. This behavior was already observed in ref. [49], where however the degrees of freedom changed throughout a non-equilibrium evolution are proportional to the size of a defect with open boundary conditions. In our case the inverse coupling $\beta$ is changed over all links, so the degrees of freedom of interest are proportional to the entire volume of the lattice. Crucially, the same scaling appears to hold also for trained SNFs: this is one of the main results of this work. We remark that the training was performed only for $n_{\mathrm{step}}=16,32,64$ and then the parameters were transferred to larger $n_{\mathrm{step}}$: yet, the same factor 2 of improvement over the NE-MCMC is always valid for all the volumes considered in this work: it appears that SNFs inherit the scaling with the volume (or, more precisely, with the degrees of freedom modified during an evolution) fully from the NE-MCMC. As it can be seen from fig. 4, the same scaling holds also for transformations in $\beta$ at coarser lattice spacings.
@@ V Conclusions and future prospects
In this work we have presented the first implementation of Stochastic Normalizing Flows for the simulation of the $\mathrm{SU}(3)$ pure gauge theory in 4 spacetime dimensions, with a particular focus on their scaling properties with the degrees of freedom involved in the transformation between the base distribution and the target distribution that we want to sample from. This architecture was tested for the case of transformations in the lattice spacing of the theory, controlled by changing the inverse coupling $\beta$ of the theory over all the links on the lattice. The scaling relation that we found was that the metrics used to evaluate the quality of the sampling do not depend separately on the length of the transformation (i.e. $n_{\mathrm{step}}$) and the effective degrees of freedom (in this case proportional to the volume $(L/a)^{4}$), but just on the ratio $n_{\mathrm{step}}/(L/a)^{4}$. This kind of scaling was already observed for NE-MCMC simulations in ref. [49] and confirmed for the purely stochastic transformation also in this case: crucially, the same scaling is inherited also by SNFs, even if they involve deterministic transformations between the out-of-equilibrium Monte Carlo steps.
@@
More generally, this work paves the way for a systematic improvement of SNFs, while retaining the desirable scaling properties found in this analysis. We envision this improvement program to develop in two different directions: the first is towards a study of optimal protocols, which have been the subject of a vast literature in stochastic thermodynamics (see ref. [62] for a review) and for which several optimization methods already exist, see for example [55, 56]. The second direction would be in the implementation of more complex and expressive gauge-equivariant layers: also in this case, recent advances in the development of Normalizing Flows for gauge theories provided generalizations of the smearing procedures in which the parameter $\rho^{(n)}$ itself is the output of a neural network [29]. We leave the study of these improvements to future work.
Concerning practical applications to critical slowing down, we would like to point out that an efficient implementation of NE-MCMC for topological freezing has already been studied for transformations in the boundary conditions, both in $\mathrm{CP}^{(N-1)}$ models in 2 dimensions (see ref. [49]) and in $\mathrm{SU}(3)$ in four dimensions, (see preliminary results in ref. [50]). This strategy involves a scaling of $n_{\mathrm{step}}$ with the size of the boundary which is switched from open to periodic boundary conditions, which in 4 dimensions amounts to a three-dimensional object, making it naturally cheaper than a transformation of $\beta$ over the whole lattice. The generalization of the SNFs presented in this work for evolutions in the boundary conditions is relatively straightforward and will be the focus of future studies.

### v2

@@ II Non-equilibrium Markov Chain Monte Carlo
A non-equilibrium Markov Chain Monte Carlo (NE-MCMC) simulation is composed of an ensemble of out-of-equilibrium evolutions from a prior distribution $q_{0}=\exp(-S_{0})/Z_{0}$ to the target distribution $p=\exp(-S)/Z$ that we wish to sample from. Each of these evolutions is generated by drawing a sequence of $n_{\mathrm{step}}$ configurations $U_{n}$ using a varying transition probability. The starting configuration $U_{0}$ of this sequence is sampled from the prior $q_{0}$ and then driven towards the target distribution $p$ through a sequence of Monte Carlo updates defined by the transition probabilities $P_{c(n)}$. These transition probabilities depend on a set of parameters $c(n)$, known as the protocol, which controls the intermediate steps of the system’s evolution, and drive the configurations sequence from the initial state toward the target distribution, i.e.,
@@
where we note that the last transition probability is fixed to be the one of the target distribution, i.e., $P_{c(n_{\mathrm{step}})}\equiv P_{p}$.
Formally, after fixing $n_{\mathrm{step}}$ and the protocol $c(n)$, we can introduce the forward $\mathcal{P}_{\mathrm{f}}$ probability distribution of each evolution $[U_{0},U_{1},\dots,U]\equiv\mathcal{U}$, which can be written as
@@
similarly, the reverse probability distribution $\mathcal{P}_{\mathrm{r}}$ takes the form of
@@
In this expression, $\Delta F=-\log Z/Z_{0}$ is the dimensionless free energy difference [^1] Across the manuscript $\Delta F$ is more precisely the free energy in units of the temperature: we remark that in a non-zero temperature setup this quantity is indeed connected to the actual free energy of the theory. We refer to ref.[44] for an in-depth description on how to compute with NE-MCMC the pressure and the equation of state in the $\mathrm{SU}(3)$ pure gauge theory. between the states described by $q_{0}$ and $p$, and $W$ represents the dimensionless work done on the system during its transformation from the initial to the final state:
@@
The computation of free energy differences between two distributions using Monte Carlo simulations is notoriously cumbersome, as it generally requires a large number of samples from intermediate distributions at equilibrium. Thus, eq. (13) represents a powerful estimator of free energy differences, as the intermediate distributions can be sampled without ever letting the system thermalize. In the context of lattice gauge theories Jarzynski’s equality has seen a wide range of applications in the last few years, see refs. [43, 44, 45, 46, 47]. Furthermore, eqs. (13) and  (11) represent the main ingredients of NE-MCMC. i.e., they provide a framework to sample problematic distributions from non-equilibrium distributions in an unbiased way; we refer to ref. [49] for a recent application on a model with frozen topology.
Note that no assumptions regarding thermalization were made during the construction of the NE-MCMC evolutions: the only exception is that if the prior distribution $q_{0}$ is sampled with a MCMC, it must be at equilibrium. While the starting and ending states of the transformations are fixed, there are in principle no restrictions on the protocol, either in the velocity of the evolution (i.e., $n_{\mathrm{step}}$) or in the functional form of $c(n)$; thus, the intermediate and final configurations $U_{n}$, with $0<n\leq n_{\mathrm{step}}$, are in principle allowed to lie arbitrarily far from equilibrium.
@@
It measures the similarity between two distributions and, in this non-equilibrium context, provides a way to study in a quantitative way the reversibility of the stochastic transformation we are employing. Large dissipation, which thermodynamically corresponds to strong irreversibility, leads to a large KL divergence; conversely, when the evolution is fully reversible forward and reverse evolutions are indistinguishable, resulting in $\tilde{D}_{\mathrm{KL}}(q_{0}\mathcal{P}_{\mathrm{f}}\|p\mathcal{P}_{\mathrm{r}})=0$. Intriguingly, since the KL divergence is positive-definite eq. (14) implies the Second Law of thermodynamics for these non-equilibrium transformations, i.e.,
@@
The KL divergence of eq. (14) represents a crucial metric to determine the reliability of NE-MCMC when sampling a finite number of evolutions. To understand why, we start by looking at the probability distribution of the last configuration $U\equiv U_{n_{\mathrm{step}}}$, i.e., the one generated by a given protocol at the end of the evolution. We can write it as:
@@
One simple and effective approach to either reduce the KL divergence of eq. (14) or to increase the ESS of eq. (19) is to increase the number of steps in the NE-MCMC protocols, i.e., going towards the $n_{\mathrm{step}}\to\infty$ limit: smoother, “quasi-static” transformations keep the thermodynamic transformations close to equilibrium, thereby maintaining vanishing dissipation. A more elaborate and potentially much more effective strategy is to optimize the functional form of the protocol $c(n)$ itself: this approach remains one of the main challenges for implementing more efficient NE-MCMC algorithms.
@@ III Stochastic Normalizing Flows
$\begin{split}W(\mathcal{U})=\sum_{n=0}^{n_{\mathrm{step}}-1}&S_{c(n+1)}\left[g_{n+1}(U_{n})\right]-S_{c(n)}\left[U_{n}\right]+\\ &-\log J_{g_{n+1}}(U_{n}).\end{split}$ (26)
@@ III.1 Gauge Equivariant Layers
The optimal stout-smearing parameters $\rho^{(n)}_{\mu\nu}(x)$ in eq. (33) have to be determined using a training procedure. In this work we simply set them to be invariant under translations and (discrete) rotations, i.e., $\rho^{(n)}_{\mu\nu}(x)=\rho^{(n)}$, for a total of 8 parameters per layer. This construction makes the Jacobian of the transformation easily tractable; for a discussion of invertibility we refer to the discussion on the link-level flow in ref. [29].
@@ IV.1 Training SNFs
The search for the optimal parameters of a SNF architecture is performed through the minimization of the KL divergence of eq. (14). This minimization procedure has a clear thermodynamic interpretation: the parameters of the intermediate stout smearing layers are tuned in such a way that the evolution becomes as reversible as possible, i.e., towards $\langle W_{\mathrm{d}}(\mathcal{U})\rangle_{\mathrm{f}}=0$. It is interesting to note that the structure of the work of eq. (26) is rather peculiar, as it can be expressed as a sum over the steps $n$, each term containing a “stochastic” contribution from the difference of the action and the “deterministic” contribution from the Jacobian of the gauge-equivariant layer. It is then very natural to ask whether the training itself can be performed step by step, i.e., backpropagating the gradients not over the whole flow, but only between two MCMC updates: in practice, one uses the loss
$S_{c(n+1)}\left[g_{n+1}(U_{n})\right]-S_{c(n)}\left[U_{n}\right]-\log J_{g_{n+1}}(U_{n})$
@@ IV.2 Scaling of NE-MCMC and SNFs
In this section we compare the performances of NE-MCMC and trained SNFs using the metrics introduced in section II, for different ensembles of transformations in the lattice spacing. In fig. 2 results for the KL divergence of eq. (14) and for the ESS of eq. (19) are presented for two volumes and for several values of $n_{\mathrm{step}}$. The dimensionless free energy $\Delta F$, needed to compute both quantities, is calculated separately for each flow using eq. (13) using the appropriate definition of the dimensionless work $W$.
@@
First of all, we observe that when $n_{\mathrm{step}}$ increases, i.e., when evolutions are performed closer to equilibrium, the KL divergence decreases for all architectures, reaching rather quickly values smaller than 1. As expected, the results indicate a convergence to 00 as well, i.e., reversibility appears to be reached asymptotically. Furthermore, if we look at the metrics for each volume separately, it appears that SNFs are roughly a factor 2 more efficient than the standard NE-MCMC: indeed, the same value of $\tilde{D}_{\mathrm{KL}}(q_{0}\mathcal{P}_{\mathrm{f}}\|p\mathcal{P}_{\mathrm{r}})$ is reached for about half the value of $n_{\mathrm{step}}$.
@@
At this point it is important to stress two relevant details. First, that for the same value of $n_{\mathrm{step}}$ SNFs also perform a stout-smearing transformation of the links for each Monte Carlo update; in our experience the increase of the computational cost is however moderate, as in our implementation they are about 25% as expensive as a full MCMC update. Second, this improvement came at a very modest cost: as explained previously, the training was performed uniquely on values of $n_{\mathrm{step}}$ which are much smaller than the ones showed in fig. 2. In practical terms, the computational effort of the training is just a small fraction of the total effort used for the sampling for larger $n_{\mathrm{step}}$. Before discussing the scaling of SNFs, it is natural to ask whether a NF (i.e., a flow without NE-MCMC steps) can reach similar performances. In this case, using the simple smearing transformation defined in eq. (30) it was not possible to create a satisfying NF, both by training the parameters from scratch or by using the ones obtained training a corresponding SNF.
Now we come to the analysis of the scaling. In fig. 3 we plotted the results for the KL divergence and the ESS as a function of the number of steps normalized by the volume in lattice units. In this way, the results of fig. 2 collapse on a unique curve, along those for $L/a=10$ and $L/a=20$: at least approximately, both metrics can be considered to be a function of $n_{\mathrm{step}}/(L/a)^{4}$. In other words, if we keep the KL divergence or the $\hat{\mathrm{ESS}}$ fixed, i.e., we fix a target for our flow-based sampling, the number of steps $n_{\mathrm{step}}$ needed to reach this target scales with the volume of the lattice. This behavior was already observed in ref. [49], where however the degrees of freedom changed throughout a non-equilibrium evolution are proportional to the size of a defect with open boundary conditions. In our case the inverse coupling $\beta$ is changed over all links, so the degrees of freedom of interest are proportional to the entire volume of the lattice. Crucially, the same scaling appears to hold also for trained SNFs: this is one of the main results of this work. We remark that the training was performed only for $n_{\mathrm{step}}=16,32,64$ and then the parameters were transferred to larger $n_{\mathrm{step}}$: yet, the same factor 2 of improvement over the NE-MCMC is always valid for all the volumes considered in this work: it appears that SNFs inherit the scaling with the volume (or, more precisely, with the degrees of freedom modified during an evolution) fully from the NE-MCMC. As it can be seen from fig. 4, the same scaling holds also for transformations in $\beta$ at coarser lattice spacings.
@@ V Conclusions and future prospects
In this work we have presented the first implementation of Stochastic Normalizing Flows for the simulation of the $\mathrm{SU}(3)$ pure gauge theory in 4 spacetime dimensions, with a particular focus on their scaling properties with the degrees of freedom involved in the transformation between the base distribution and the target distribution that we want to sample from. This architecture was tested for the case of transformations in the lattice spacing of the theory, controlled by changing the inverse coupling $\beta$ of the theory over all the links on the lattice. The scaling relation that we found was that the metrics used to evaluate the quality of the sampling do not depend separately on the length of the transformation (i.e., $n_{\mathrm{step}}$) and the effective degrees of freedom (in this case proportional to the volume $(L/a)^{4}$), but just on the ratio $n_{\mathrm{step}}/(L/a)^{4}$. This kind of scaling was already observed for NE-MCMC simulations in ref. [49] and confirmed for the purely stochastic transformation also in this case: crucially, the same scaling is inherited also by SNFs, even if they involve deterministic transformations between the out-of-equilibrium Monte Carlo steps.
@@
More generally, this work paves the way for a systematic improvement of SNFs, while retaining the desirable scaling properties found in this analysis. We envision this improvement program to develop in two different directions: the first is towards a study of optimal protocols, which have been the subject of a vast literature in stochastic thermodynamics (see ref. [62] for a review) and for which several optimization methods already exist, see for example [55, 56]. The second direction would be in the implementation of more complex and expressive gauge-equivariant layers: also in this case, recent advances in the development of Normalizing Flows for gauge theories provided generalizations of the smearing procedures in which the parameter $\rho^{(n)}$ itself is the output of a neural network [29]. Although we leave the study of these improvements to future work, we wish to remark that a more comprehensive study of the relationship between gauge-equivariant coupling layers and the non-equilibrium Monte Carlo updates would be extremely valuable. An interesting aspect to examine in depth would be the performance of a SNF with a relatively small number of NE-MCMC steps but enhanced by more complex and expressive gauge-equivariant coupling layers. We expect that such an architecture would involve improved efficiency, as the flow would be less constrained by the protocol; conversely, it would likely require more significant training costs as well. Furthermore, it would be interesting to explore if (and when) the scaling with the degrees of freedom induced by the NE-MCMC breaks down, as the complexity of the network increases and/or the number of Monte Carlo steps decreases. We point out that a comparison between NFs and SNFs has been carried out in [24], albeit in a simpler setting.
Concerning practical applications to critical slowing down, we would like to point out that an efficient implementation of NE-MCMC for topological freezing has already been studied for transformations in the boundary conditions, both in $\mathrm{CP}^{(N-1)}$ models in 2 dimensions (see ref. [49]) and in $\mathrm{SU}(3)$ in four dimensions (see preliminary results in ref. [50]). This strategy involves a scaling of $n_{\mathrm{step}}$ with the size of the boundary which is switched from open to periodic boundary conditions, which in 4 dimensions amounts to a three-dimensional object, making it naturally cheaper than a transformation of $\beta$ over the whole lattice. The generalization of the SNFs presented in this work for evolutions in the boundary conditions is relatively straightforward and will be the focus of future studies.